A 5.0kg ball is projected up along a plane that is inclined 20 degrees with the horizontal. What is its potential emergy when it moved 2.3m up along the plane?​


1
Expert's answer
2021-09-29T09:54:12-0400

The potential energy

"E_p=mgh=mgl\sin\theta\\
=\rm5.0\: kg*9.8\: m/s^2*2.3\: m*\sin20^{\circ}=38.5\: J"